# bruhat.bib

### v1

@@
[…]
bruhat.bib
\@setabstract
Explicit formula for the $(\mathrm{GL}_{2},\mathrm{GL}_{2})$ theta lift via Bruhat decomposition
Peter Xu
@@ 1.1. Related work and future directions
A similar computation using complexes to compoute equivariant cohomology was completed in [KS, §3], in a more general setting using a generalized Kronecker-Eisenstein series. Our calculation can be considered as a more explicated form of the calculation there, for the self-product of the universal elliptic curve with trivial coefficients, except that we do not rely on any stabilization or smoothing coming from a degree-zero torsion cycle, and we construct also a cocycle for a larger group with no geometric action on the fibers ($\mathrm{GL}_{2}(\mathbb{Q})$ instead of $\mathrm{GL}_{2}(\mathbb{Z})$). We restrict to this setting both because it is particularly arithmetically rich, and allows for a considerably simpler and less technical exposition. Futhermore, by eschewing smoothing, the resulting formulas are more transparently related to classical objects such as classical holomorphic Eisenstein series or the period formulas of Siegel (as in [X4]), and corresponds directly to the theta kernel of integration of [BCG]. Furthermore, our formulas can be specialized (with cohomological meaning) at arbitrary non-zero torsion points, thanks to Theorem 3.5.
As we noted in our previous article [X4] in a simpler setting, our same methods should produce analogous formulas with twisted coefficients corresponding to higher-weight automorphic sheaves; the resulting formulas will involve progressively more complicated combinations of higher-weight specializations of Kronecker-Eisenstein series. Another natural generalization would be to use the more complicated decomposition of parabolic subgroups for $\mathrm{GL}_{n}$ and $n$-fold products of elliptic curves; in this setting, our unsmoothed formulas grow in complexity very quickly with $n$. Furthermore, the direct analogue of Theorem 3.5 does not hold for $n>2$, so the ability to specialize at torsion sections is more limited. Nevertheless, in the smoothed setting, things are much simpler and can be captured by symbols defined linear algebraically; we follow this approach for elliptic schemes in [X3].
@@ 4.1. Hecke equivariance
Pulling back this equality by $x$ (since this commutes with the pullback action of $\Delta$!), we obtain:
@@ 4.2. CM elliptic curves
This imaginary quadratic cocycle can be viewed as being “for the dual pair $(\mathrm{GL}_{2}(K),\mathrm{GL}_{1}(K))$,” and is approached analytically in the work [BCGIm]. Similarly interesting cocycles can be constructed analogously also for products of two isogenous CM elliptic curves corresponding to different elements of the class group $\text{Cl}(K)$, but we do not go into this year.

### v2

@@
Abstract.
Using combinations of weight-$1$ and weight-$2$ of Kronecker-Eisenstein series to construct currents in the distributional de Rham complex of a squared elliptic curve, we find a simple explicit formula for the type II $(\mathrm{GL}_{2},\mathrm{GL}_{2})$ theta lift without smoothing, analogous to the classical formula of Siegel for periods of Eisenstein series. For $K$ a CM field, the same technique applies without change to obtain an analogous formula for the $(\mathrm{GL}_{2}(K),K^{\times})$ theta correspondence.
Utilisant des combinaisons des séries et courants Kronecker-Eisenstein de poids $1$ et $2$ dans des complexes de Rham d’une courbe elliptique carrée, nous trouvons une formule simple explicite pour la correspondance thêta de type II pour $(\mathrm{GL}_{2},\mathrm{GL}_{2})$ sans stabilisation; cela est un analogue de la formule classique de Siegel pour les périodes des séries d’Eisenstein. Pour $K$ un corps avec CM, la même technique s’applique pareil pour obtenir une telle formule pour la correspondance $(\mathrm{GL}_{2}(K),K^{\times})$.
@@ 1.1. Related work and future directions
A similar computation using complexes to compoute equivariant cohomology was completed in [KS, §3], in a more general setting using a generalized Kronecker-Eisenstein series. Our calculation can be considered as a more explicated form of the calculation there, for the self-product of the universal elliptic curve with trivial coefficients, except that:
  1. (1)
    we treat the entire action of $\mathrm{GL}_{2}$ at once, instead of just working with the equivariance under one subtorus at a time;
  2. (2)
    we construct also a cocycle for a larger group with no geometric action on the fibers ($\mathrm{GL}_{2}(\mathbb{Q})$ instead of $\mathrm{GL}_{2}(\mathbb{Z})$, or $\mathrm{GL}_{2}(K)$ in the CM case);
  3. (3)
    we do not rely on any stabilization or smoothing coming from a degree-zero torsion cycle.
We restrict to this setting both because it is particularly arithmetically rich, and allows for a considerably simpler and less technical exposition. Futhermore, by eschewing smoothing, the resulting formulas are more transparently related to classical objects such as classical holomorphic Eisenstein series or the period formulas of Siegel (as in [X4]), and corresponds directly to the theta kernel of integration of [BCG]. Furthermore, our formulas can be specialized (with cohomological meaning) at arbitrary non-zero torsion points, thanks to Theorem 3.5.
As we noted in our previous article [X4] in a simpler setting, our same methods should produce analogous formulas with twisted coefficients corresponding to higher-weight automorphic sheaves; the resulting formulas will involve progressively more complicated combinations of higher-weight specializations of Kronecker-Eisenstein series. Another natural generalization would be to use the more complicated decomposition of parabolic subgroups for $\mathrm{GL}_{n}$ and $n$-fold products of elliptic curves; in this setting, our unsmoothed formulas grow in complexity very quickly with $n$, and it is not clear to us if there exists a symbol formalism that can help one understand this complexity. Furthermore, the direct analogue of Theorem 3.5 does not hold for $n>2$, so the ability to specialize at torsion sections is more limited. Nevertheless, in the smoothed setting, things are much simpler and can be captured by symbols coming from elementary linear algebra; we follow this approach for elliptic schemes in [X3].
1.2. Acknowledgements
I would like to thank first and foremost Nicolas Bergeron, to whose ideas I owe all of my work in this subject area; and who very kindly accommodated me in Paris in February 2023, listening to and helping to develop my ideas even with the limited energy he had at the time. This article is dedicated in his memory. I would also like to thank Romyar Sharifi, Marti Roset Julia, and Pierre Charollois for helpful consultations during the writing process.
@@ 4.1. Hecke equivariance
Pulling back this equality by $x$ (since this commutes with the pullback action of $\Delta$), we obtain:
@@
Using this Hecke equivariance, one can obtain explicit spectral decompositions of the $(\mathrm{GL}_{2},\mathrm{GL}_{2})$-theta lift, by using the Rankin-Selberg formula for the inner product of an eigenform with the $E_{1}E_{1}$ terms in (3.21). However, we will pursue this via a more systematic approach in future work, so do not go into it here.
@@ 4.2. CM elliptic curves
This imaginary quadratic cocycle can be viewed as being “for the dual pair $(\mathrm{GL}_{2}(K),\mathrm{GL}_{1}(K))$,” and is approached analytically in the work [BCGIm]. In particular, our formula above gives a simple expression in terms of Eisenstein-Kronecker numbers of the weight-$(0,0)$ cocycle denoted $\Phi^{0,0}$ in loc. cit., and therefore also the values of Hecke $L$-functions associated to weight-00 characters for the field $K$, as in [BCGIm, Theorem 1.2]. As in the $\mathrm{GL}_{2}(\mathbb{Q})$ case, by employing twisted versions of our complexes and taking different weight Eisenstein-Kronecker series, it is possible to obtain analogous formulas for the more general cocycles $\Phi^{p,q}$, which could be an interesting direction of future work.